# A progress on the binary Goldbach conjecture

### v8

@@ Abstract.
In this paper we develop the method of circle of partitions and associated statistics. As an application we prove conditionally the binary Goldbach conjecture. We develop series of steps to prove the binary Goldbach conjecture in full. We end the paper by proving the binary Goldbach conjecture for all sufficiently large even numbers.
@@ 1. Introduction
The Goldbach conjecture dates from $1742$ out of the correspondence between the Swiss mathematician Leonard Euler and the German mathematician Christian Goldbach. The problem has two folds, namely the binary case and the ternary case. The binary case ask if every even number $\geq 6$ can be written as a sum of two primes, where as the ternary case ask if every odd number $\geq 7$ can be written as a sum of three prime numbers. The ternary case has, however, been solved quite recently in the preprint [2] culminating several works. Though the binary problem remains unsolved as of now there has been substantive progress as well as on its variants. The first milestone in this direction can be found in (see [5]), where it is shown that every even number can be written as the sum of at most $C$ primes, where $C$ is an effectively computable constant. In the early twentieth century, G.H Hardy and J.E Littlewood assuming the Generalized Riemann hypothesis (see [8]), showed that the number of even numbers $\leq X$ and violating the binary Goldbach conjecture is much less than $X^{\frac{1}{2}+c}$, where $c$ is a small positive constant. Jing-run Chen [3], using the methods of sieve theory, showed that every even number can either be written as a sum of two prime numbers or a prime number and a number which is a product of two primes. It also known that almost all even numbers can be written as the sum of two prime numbers, in the sense that the density of even numbers representable in this manner is one [7], [6]. It is also known that there exist a constant $K$ such that every even number can be written as the sum of two prime numbers and at most $K$ powers of two, where we can take $K=13$ [4].
In [10] we have developed a method which we feel might be a valuable resource and a recipe for studying problems concerning partition of numbers in specified subsets of $\mathbb{N}$. The method is very elementary in nature and has parallels with configurations of points on the geometric circle.
Let us suppose that for any $n\in\mathbb{N}$ we can write $n=u+v$ where $u,v\in\mathbb{M}\subset\mathbb{N}$ then the new method associate each of this summands to points on the circle generated in a certain manner by $n>2$ and a line joining any such associated points on the circle. This geometric correspondence turns out to useful in our development, as the results obtained in this setting are then transformed back to results concerning the partition of integers.
@@
Let $n\in\mathbb{N}$ and $\mathbb{M}\subseteq\mathbb{N}$. We denote with
$\mathcal{C}(n,\mathbb{M})=\left\{[x]\mid x,y\in\mathbb{M},n=x+y\right\}$
@@
the Circle of Partition generated by $n$ with respect to the subset $\mathbb{M}$. We will abbreviate this in the further text as CoP. We call members of $\mathcal{C}(n,\mathbb{M})$ as points and denote them by $[x]$. For the special case $\mathbb{M}=\mathbb{N}$ we denote the CoP shortly as $\mathcal{C}(n)$. We denote with $\|[x]\|:=x$ the weight of the point $[x]$ and correspondingly the weight set of points in the CoP $\mathcal{C}(n,\mathbb{M})$ as $\|\mathcal{C}(n,\mathbb{M})\|$. Obviously holds
@@
We denote $\mathbb{L}_{[x],[y]}$ as an axis of the CoP $\mathcal{C}(n,\mathbb{M})$ if and only if $x+y=n$. We say the axis point $[y]$ is an axis partner of the axis point $[x]$ and vice versa. We do not distinguish between $\mathbb{L}_{[x],[y]}$ and $\mathbb{L}_{[y],[x]}$, since it is essentially the same axis. The point $[x]\in\mathcal{C}(n,\mathbb{M})$ such that $2x=n$ is the center of the CoP. If it exists then, we call it as a degenerated axis $\mathbb{L}_{[x]}$ in comparison to the real axes $\mathbb{L}_{[x],[y]}$. We denote the assignment of an axis $\mathbb{L}_{[x],[y]}$ to a CoP $\mathcal{C}(n,\mathbb{M})$ as
$\mathbb{L}_{[x],[y]}~{}\hat{\in}~{}\mathcal{C}(n,\mathbb{M})\mbox{ which means}[x],[y]\in\mathcal{C}(n,\mathbb{M})\mbox{ with }x+y=n.$
@@
In the following sequel, we consider only real axes. That is, axes of the form $\mathbb{L}_{[x],[y]}$ such that $x\neq y$. Therefore, we abstain from the attribute real in the sequel. The introduction of the notation $||[x]||:=x$ denoting the weight of the point $[x]$ is not superfluous, as there are situations where the point is subject to motions as rotation and dilation or possibly flipping and using this notation becomes necessary. On a more advisory note, we have found the approach in the paper the ideal and possibly the perfect language to study additive problems of this kind. The structure under study has various connections with the structure of the geometric circle and much of our intuition has been borrowed from this setting.
@@
Each axis is uniquely determined by points $[x]\in\mathcal{C}(n,\mathbb{M})$.
@@
Let $\mathbb{L}_{[x],[y]}$ be an axis of the CoP $\mathcal{C}(n,\mathbb{M})$. Suppose as well that $\mathbb{L}_{[x],[z]}$ is also an axis with $z\neq y$. Then we must have $n=x+y=x+z$ and therefore $y=z$. This cannot be and the claim follows immediately. ∎
@@
Let $[x]\in\mathcal{C}(n,\mathbb{M})$ be a point without an axis partner being not the center of the CoP. Then holds for every point $[y]\neq[x]$ except the center
@@
This is impossible, since each point in a CoP must have an axes partner. Due to Proposition 1.2 the case of more than one axis partners is impossible. This completes the proof. ∎
Notation
We denote by
$\mathbb{N}_{n}=\left\{m\in\mathbb{N}\mid~{}m\leq n\right\}$ (1.1)
the sequence of the first $n$ natural numbers. We denote the assignment of an axis $\mathbb{L}_{[x],[y]}$ resp. $\mathbb{L}_{[x]}$ to a CoP $\mathcal{C}(n,\mathbb{M})$ as
$\mathbb{L}_{[x],[y]}~{}\hat{\in}~{}\mathcal{C}(n,\mathbb{M})\mbox{ which means }[x],[y]\in\mathcal{C}(n,\mathbb{M})\mbox{ and }x+y=n\mbox{ resp.}$.
$\mathbb{L}_{[x]}~{}\hat{\in}~{}\mathcal{C}(n,\mathbb{M})\mbox{ which means }[x]\in\mathcal{C}(n,\mathbb{M})\mbox{ and }2x=n$
and the number of real axes of a CoP as
@@
Obviously holds
$\nu(n,\mathbb{M})=\left\lfloor\frac{k}{2}\right\rfloor,\mbox{ if }|\mathcal{C}(n,\mathbb{M})|=k.$
@@ 2. The Density of Points on the Circle of Partition
In this section we introduce the notion of density of points on CoP $\mathcal{C}(n,\mathbb{M})$ for $\mathbb{M}\subseteq\mathbb{N}$. We launch the following language in that regard. We consider in this section only real axes. Hence, we refrain from the use of the attribute real in this section.
@@
Let be $\mathbb{H}\subset\mathbb{N}$. Then the limits
@@
denotes the density of $\mathbb{H}$ if it exists.
@@
Let $\mathcal{C}(n,\mathbb{M})$ be CoP with $\mathbb{M}\subset\mathbb{N}$ and $n\in\mathbb{N}$. Suppose $\mathbb{H}\subset\mathbb{M}$ then by the density of points $[x]\in\mathcal{C}(n,\mathbb{M})$ such that $x\in\mathbb{H}$, denoted $\mathcal{D}(\mathbb{H}_{\mathcal{C}(\infty,\mathbb{M})})$, we mean the limit
$\mathcal{D}\left(\mathbb{H}_{\mathcal{C}(\infty,\mathbb{M})}\right)=\lim\limits_{n\longrightarrow\infty}\frac{\#\{\mathbb{L}_{[x],[y]}~{}\hat{\in}~{}\mathcal{C}(n,\mathbb{M})|~{}\{x,y\}\cap\mathbb{H}\neq\emptyset\}}{\nu(n,\mathbb{M})}$
if it exists.
The notion of the density of points as espoused in Definition 2.2 provides a passage between the density of the corresponding weight set of points. This possibility renders this type of density as a black box in studying problems concerning partition of numbers into specialized sequences taking into consideration their density.
@@
Let $\mathcal{C}(n)$ with $n\in\mathbb{N}$ be a CoP and $\mathbb{H}\subset\mathbb{N}$. Then the following inequality holds
@@
The upper bound is obtained from a configuration where no two points $[x],[y]\in\mathcal{C}(n)$ such that $x,y\in\mathbb{H}$ lie on the same axis of the CoP. That is, by the uniqueness of the axes of CoPs with $\nu(n,\mathbb{H})=0$, we can write
#$\#\left\{\mathbb{L}_{[x],[y]}\in\mathcal{C}(n)|~{}\{x,y\}\cap\mathbb{H}\neq\emptyset\right\}$ $=\nu(n,\mathbb{H})+\#\left\{\mathbb{L}_{[x],[y]}\in\mathcal{C}(n)|~{}x\in\mathbb{H},~{}y\in\mathbb{N}\setminus\mathbb{H}\right\}$
=#$=\#\left\{\mathbb{L}_{[x],[y]}\in\mathcal{C}(n)|~{}x\in\mathbb{H},~{}y\in\mathbb{N}\setminus\mathbb{H}\right\}$
@@
The lower bound however follows from a configuration where any two points $[x],[y]\in\mathcal{C}(n)$ with $x,y\in\mathbb{H}$ are joined by an axis of the CoP. That is, by the uniqueness of the axis of CoPs with #$\#\left\{\mathbb{L}_{[x],[y]}\in\mathcal{C}(n)|~{}x\in\mathbb{H},~{}y\in\mathbb{N}\setminus\mathbb{H}\right\}=0$, then we can write
#$\#\left\{\mathbb{L}_{[x],[y]}\in\mathcal{C}(n)|~{}\{x,y\}\cap\mathbb{H}\neq\emptyset\right\}$ $=\nu(n,\mathbb{H})$
@@
Let $\mathbb{H}\subset\mathbb{N}$ and $\mathcal{D}(\mathbb{H}_{\mathcal{C}(\infty)})$ be the density of the corresponding points with weight set $\mathbb{H}$. Then the following properties hold:
@@
It is easy to see that the first part of Property $(i)$ and $(iii)$ are both easy consequences of the definition of density of points on the CoP $\mathcal{C}(n)$ and Proposition 2.3. We establish the second part of property $(i)$ and Property $(ii)$, which is the less obvious case. We observe by the uniqueness of the axes of CoPs that we can write
11 11 $=\lim\limits_{n\longrightarrow\infty}\frac{\nu(n,\mathbb{N})}{\nu(n,\mathbb{N})}$
@@
Next we transfer the notion of the density of a sequence to the density of corresponding points on the CoP $\mathcal{C}(n)$. This notion will play a crucial role in our latter developments.
@@
Let $\epsilon\in(0,1]$ and $\mathbb{H}$ be a sequence with $\mathbb{H}\subset\mathbb{N}$ and $\mathcal{C}(n)$ be a CoP. If $\mathcal{D}\left(\mathbb{H}\right)\geq\epsilon$ then $\mathcal{D}\left(\mathbb{H}_{\mathcal{C}(\infty)}\right)\geq\epsilon$.
@@
2.1. Application of Density of Points to Partitions
In this subsection we explore the connection between the notion of density of points in a typical CoP to the possibility of partitioning number into certain sequences. This method tends to work very efficiently for sets of integers having a positive density.

### v9

@@ Abstract.
In this paper, we develop the method of the circle of partitions and associated statistics. As an application, we conditionally prove the binary Goldbach conjecture. We develop a series of steps to prove the binary Goldbach conjecture in full. We end the paper by proving the binary Goldbach conjecture for all sufficiently large even numbers.
@@ 1. Introduction
The Goldbach conjecture is one of the oldest and most celebrated problems in additive number theory. Its modern statement arose in the correspondence between Leonhard Euler and Christian Goldbach in 1742: every even integer greater than $2$ is the sum of two primes. Over the last century the problem has inspired a wide network of ideas and techniques—ranging from combinatorial and sieve methods to deep analytic tools—and has produced a substantial body of partial results and heuristics that illuminate the landscape around the conjecture. Early milestones include the Shnirel’man combinatorial breakthrough showing that every sufficiently large integer is a sum of a bounded number of primes [4], and the foundational contributions of Hardy and Littlewood which, under standard analytic hypotheses, elucidated the expected asymptotic distribution of prime representations [7]. Subsequent work of Estermann and Chudakov established that almost all even integers are representable as a sum of two primes [5, 6]; sieve methods produced near-miss results such as the Chen theorem on representations by a prime plus a product of at most two primes [2]; and refinements that combine sieves with exponential-sum techniques and explicit prime estimates have led to strong conditional and unconditional density results (see, e.g., [3, 8]). More recently the ternary Goldbach conjecture (every odd $n\geq 7$ is a sum of three primes) was resolved by Helfgott [1], closing a long-standing chapter while leaving the binary case still open.
In this paper, we develop and refine a geometric-combinatorial framework–the method of circles of partition (CoP) and its complex extension (cCoP)–and show how this framework can be used to transfer density information about suitable auxiliary subsets of the integers into concrete additive statements. The central idea is straightforward to describe: encode representations $n=u+v$ with $u,v$ drawn from a base set $\mathbb{M}\subset\mathbb{N}$ as pairs of “points” on a circle attached to the generator $n$. Axes joining complementary points capture the additive structure, and one studies statistics of these axes (their frequency, density, and configurations) to deduce partition properties of $n$. When the base set has positive natural density this transfer is immediate and powerful; the delicate part–and the heart of this paper–is to design auxiliary base-sets and analytic constructions that are sufficiently large and flexible to contain many primes while still admitting geometric control.
The main technical novelties of the present work are threefold. First, in Section 2, we introduce a refined notion of point-density on CoPs which mediates between the usual asymptotic density of a set $\mathbb{H}\subset\mathbb{N}$ and the fraction of axes in $\mathcal{C}(n,\mathcal{M})$ that meet $\mathbb{H}$. This viewpoint yields simple, transparent inequalities (see Proposition 3.3) that isolate the combinatorial constraints any auxiliary base-set must satisfy in order to force prime–prime axes. Second, we extend the CoP formalism to the complex plane (the cCoP), imposing a “circle condition” that ties real and imaginary parts and allows for additional symmetries (conjugate axes, embedding circles) which are useful in constructing intermediate generators. The complex viewpoint is not intended to introduce new analytic difficulties but rather to expose an ordering principle among axes that is invisible in the purely real model. Third, building on these two constructions we formulate and prove a squeeze principle (Lemmas 4.2 and 4.3) and a generalized multivariate version (Lemma 5.4). These principles are combinatorial-geometric mechanisms that allow one to “trap” an intermediate generator between two nonempty CoPs (or cCoPs) and thereby produce new axes whose residents belong to a prescribed subset. The squeeze principle plays the role of a local covering or exhaustion lemma in our argument but has the advantage of being elementary, transparent, and adaptable to multivariate partitions.
Using the framework above, we obtain two kinds of results. On the one hand, we give conditional partition statements that reduce the binary Goldbach problem to the construction of an auxiliary base set $\mathbb{B}$ for which primes occupy a proportion >1/2>1/2 of the weight-set of $\mathcal{C}(n,\mathbb{B})$ (Theorem 3.8). On the other hand, by combining the squeeze principle with explicit prime estimates (notably explicit prime-density and short-interval results such as those in [8]), we obtain an asymptotic form of the binary Goldbach conjecture: every sufficiently large even integer is a sum of two primes (Theorem 4.10). The asymptotic theorem presented here is in the spirit of earlier conditional or density-based approaches, but it is based on new combinatorial geometry and a systematic exploitation of CoP/cCoP configurations and thus provides a conceptual alternative route that complements classical sieve and circle-method strategies (compare [2, 3]).
1.1. Organization of the paper.
The paper is organized as follows. In Section 2, we introduce the circle-of-partition formalism, define axes and point-density, and record basic combinatorial properties and inequalities that connect set-density to axis-density. Section 3 develops the density-to-partition machinery and proves Theorem 3.6, illustrating the method on dense subsets (the square-free integers give a short, illuminating corollary). In Section 4, we introduce the complex circles of partition (cCoPs), develop the conjugacy and embedding-circle machinery, and prove the squeeze principle and its specializations (Lemmas 4.2 and 4.3). Section 4.1 contains the main applications: the conditional reduction (Theorem 3.8), the asymptotic squeeze arguments based on explicit prime estimates (Lemma 4.9), and the asymptotic binary Goldbach theorem (Theorem 4.10). Finally, Section 5 generalizes the squeeze principle to multivariate circles and explains how the same geometric ideas extend to $h$-fold partitions; the resulting partition law clarifies when finitely many local axis-configurations imply global partitioning statements for all large integers.
Throughout the paper, we compare and contrast the CoP viewpoint with classical approaches: the CoP formalism is deliberately elementary and combinatorial, yet it has a natural interface with analytic inputs (prime counting and explicit short-interval results) when those are required. This work builds on the preliminary CoP ideas developed in [9] and on earlier asymptotic investigations [10], and it makes explicit how geometric encoding can be combined with explicit prime bounds (e.g. [8]) to reach asymptotic Goldbach-type conclusions.
Acknowledgments. The author thanks colleagues and referees who provided helpful comments on earlier drafts and acknowledges that parts of the CoP formalism were inspired by classical combinatorial ideas in additive number theory (see, e.g., [4, 7, 2, 5, 6, 3, 1, 8]).
2. The circle of partitions
In this section, we introduce the concept of the circle of partition.
Definition 2.1.
@@ 2. The circle of partitions
Let $n\in\mathbb{N}$ and $\mathbb{M}\subseteq\mathbb{N}$. We denote the Circle of Partition generated by $n$ with respect to the subset $\mathbb{M}$ by
$\mathcal{C}(n,\mathbb{M})=\left\{[x]~|~x,y\in\mathbb{M},n=x+y\right\}$
@@
In the following, we will abbreviate this as CoP. We call members of $\mathcal{C}(n,\mathbb{M})$ points and denote them by $[x]$. For the special case $\mathbb{M}=\mathbb{N}$, we denote the CoP in a short form as $\mathcal{C}(n)$. We denote the weight of the point $[x]$ by $\|[x]\|:=x$ and, correspondingly, the weight set of points in the CoP $\mathcal{C}(n,\mathbb{M})$ by $\|\mathcal{C}(n,\mathbb{M})\|$. Clearly, we have
@@
We denote an axis of the CoP $\mathcal{C}(n,\mathbb{M})$ by $\mathbb{L}_{[x],[y]}$ if and only if $x+y=n$. We say that axis point $[y]$ is an axis partner of the axis point $[x]$ and vice versa. We do not distinguish between the axes $\mathbb{L}_{[x],[y]}$ and $\mathbb{L}_{[y],[x]}$, because it is essentially the same axis. The point $x]\in\mathcal{C}(n,\mathbb{M})$ such that $2x=n$ is the center of the CoP. If it exists, then we call it a degenerated axis and denote by $\mathbb{L}_{[x]}$ in comparison to the real axes which we denote by $\mathbb{L}_{[x],[y]}$. We denote the assignment of an axis $\mathbb{L}_{[x],[y]}$ to a CoP $\mathcal{C}(n,\mathbb{M})$ by
$\mathbb{L}_{[x],[y]}~\hat{\in}~\mathcal{C}(n,\mathbb{M})$
which means
$x],[y]\in\mathcal{C}(n,\mathbb{M})\quad\text{with}\quad x+y=n.$
@@
In the following, we consider only real axes. That is, axes of the form $\mathbb{L}_{[x],[y]}$ such that $x\neq y$. Therefore, we refrain from the attribute real in the sequel. The introduction of the notation [x]:=x||[x]||:=x denoting the weight of the point $[x]$ is not superfluous, because there are situations where the point is subject to motions by rotation and dilation or possibly flipping and using this notation becomes necessary. On a more advisory note, we have found the approach in the paper the ideal and possibly the perfect language to study additive problems of this type. The structure under study has various connections with the structure of the geometric circle and much of our intuition has been borrowed from this setting.
@@
Each axis is uniquely determined by points $x]\in\mathcal{C}(n,\mathbb{M})$.
@@
Let $\mathbb{L}_{[x],[y]}$ be an axis of the CoP $\mathcal{C}(n,\mathbb{M})$. Suppose that $\mathbb{L}_{[x],[z]}$ is also an axis with $z\neq y$. We have $n=x+y=x+z$ and therefore $y=z$. ∎
@@
Let $x]\in\mathcal{C}(n,\mathbb{M})$ be a point which is not the center of the CoP and without an axis partner . It implies that for every point $y]\neq[x]$ except the center
@@
This is impossible, since each point in a CoP must have an axes partner. Due to Proposition 2.3 the case of more than one axis partners is impossible. ∎
2.1. Notations
We denote the sequence of the first $n$ positive integers by
$\mathbb{N}_{n}=\left\{m\in\mathbb{N}~|~m\leq n\right\}.$
We denote the assignment of an axis $\mathbb{L}_{[x],[y]}$ (resp. $\mathbb{L}_{[x]}$) to a CoP $\mathcal{C}(n,\mathbb{M})$ by
$\mathbb{L}_{[x],[y]}~\hat{\in}~\mathcal{C}(n,\mathbb{M})$
which means
$x],[y]\in\mathcal{C}(n,\mathbb{M})\quad\text{and}\quad x+y=n$
respectively
$\mathbb{L}_{[x]}~\hat{\in}~\mathcal{C}(n,\mathbb{M})$
which means
$x]\in\mathcal{C}(n,\mathbb{M})\quad\text{and}\quad 2x=n$
and the number of real axes of a CoP by
@@ 2.1. Notations
It is observed that
$\nu(n,\mathbb{M})=\left\lfloor\frac{k}{2}\right\rfloor\quad\text{if}\quad|\mathcal{C}(n,\mathbb{M})|=k.$
@@ 3. The Density of Points on the Circle of Partition
In this section, we introduce the notion of density of points on CoP $\mathcal{C}(n,\mathbb{M})$ for $\mathbb{M}\subseteq\mathbb{N}$.
@@
Let $\mathbb{H}\subset\mathbb{N}$. The limits
@@
is the density of $\mathbb{H}$ if it exists and is finite.
@@
Let $\mathcal{C}(n,\mathbb{M})$ be CoP with $\mathbb{M}\subset\mathbb{N}$ and $n\in\mathbb{N}$ with $\mathbb{H}\subset\mathbb{M}$. The density of points $x]\in\mathcal{C}(n,\mathbb{M})$ such that $x\in\mathbb{H}$, denoted by $\mathcal{D}(\mathbb{H}_{\mathcal{C}(\infty,\mathbb{M})})$, is the limit
$\mathcal{D}\left(\mathbb{H}_{\mathcal{C}(\infty,\mathbb{M})}\right)=\lim\limits_{n\longrightarrow\infty}\frac{\#\{\mathbb{L}_{[x],[y]}~\hat{\in}~\mathcal{C}(n,\mathbb{M})~|~\{x,y\}\cap\mathbb{H}\neq\emptyset\}}{\nu(n,\mathbb{M})}$
if it exists and is finite.
The notion of the density of points as espoused in the definition 3.2 allows a passage between the density of the corresponding weight set of points. This possibility makes this type of density a black box for the study of problems concerning the partition of numbers into specialized sequences taking into consideration their density.
@@
Let $\mathcal{C}(n)$ with $n\in\mathbb{N}$ be a CoP and $\mathbb{H}\subset\mathbb{N}$. The following inequality holds:
@@
The upper bound is obtained from a configuration in which no two points $x],[y]\in\mathcal{C}(n)$ such that $x,y\in\mathbb{H}$ lie on the same axis of the CoP. That is, by the uniqueness of the axes of CoPs with $\nu(n,\mathbb{H})=0$, we can write
$\#\left\{\mathbb{L}_{[x],[y]}\in\mathcal{C}(n)|~\{x,y\}\cap\mathbb{H}\neq\emptyset\right\}$ $=\nu(n,\mathbb{H})+\#\left\{\mathbb{L}_{[x],[y]}\in\mathcal{C}(n)|~x\in\mathbb{H},~y\in\mathbb{N}\setminus\mathbb{H}\right\}$
$=\#\left\{\mathbb{L}_{[x],[y]}\in\mathcal{C}(n)|~x\in\mathbb{H},~y\in\mathbb{N}\setminus\mathbb{H}\right\}$
@@
However, the lower bound follows from a configuration where any two points $x],[y]\in\mathcal{C}(n)$ with $x,y\in\mathbb{H}$ are joined by an axis of the CoP. That is, by the uniqueness of the axis of CoPs with
$\left\{\mathbb{L}_{[x],[y]}\in\mathcal{C}(n)~|~x\in\mathbb{H},~y\in\mathbb{N}\setminus\mathbb{H}\right\}=0$
we can write
$\#\left\{\mathbb{L}_{[x],[y]}\in\mathcal{C}(n)~|~\{x,y\}\cap\mathbb{H}\neq\emptyset\right\}$ $=\nu(n,\mathbb{H})$
@@
Let $\mathbb{H}\subset\mathbb{N}$ and $\mathcal{D}(\mathbb{H}_{\mathcal{C}(\infty)})$ be the density of the corresponding points with weight set $\mathbb{H}$. The following properties hold:
@@
It observed that the first part of Property $(i)$ and $(iii)$ are both easy consequences of the definition of density of points on the CoP $\mathcal{C}(n)$ and Proposition 3.3. We establish the second part of property $(i)$ and Property $(ii)$, which is the less obvious case. We observe, by the uniqueness of the axes of CoPs, that we can write
$1$ $=\lim\limits_{n\longrightarrow\infty}\frac{\nu(n,\mathbb{N})}{\nu(n,\mathbb{N})}$
@@
Here, we transfer information of the density of a sequence to the density of corresponding points on the CoP $\mathcal{C}(n)$. This approach will play an important role in our latter developments.
@@
Let $\epsilon\in(0,1]$ and $\mathbb{H}\subset\mathbb{N}$ and $\mathcal{C}(n)$ be a CoP. If $\mathcal{D}\left(\mathbb{H}\right)\geq\epsilon$, then $\mathcal{D}\left(\mathbb{H}_{\mathcal{C}(\infty)}\right)\geq\epsilon$.
@@
3.1. Application of the density of Points to Partitions
Here, we apply the notion of the density of points in a typical CoP to write an integer as the sum of two integers that belong a specific subset of the integers. This method works efficiently for sets of integers having a positive density.